# Large-scale semi-organized rolls in a sheared convective turbulence: Mean-field simulations

### v1

@@ IV Results of mean-field numerical simulations
In this section, we discuss results of mean-field numerical simulations for three-dimensional problem. We solve numerically Eqs. (4)–(5) for the periodic boundary conditions in the horizontal $xy$ plane. The boundary conditions for the potential temperature in the vertical direction are $\tilde{\Theta}(t,z=0)=\tilde{\Theta}(t,z=1)=0$. We use the stress-free and no-slip boundary conditions for the velocity field in the vertical direction (along the $z$ axis). The stress-free boundary conditions imply,
@@
We use the ANSYS FLUENT code (version 19.2) in the $3D$ box $L_{x}=5$ and $L_{y}=L_{z}=1$, which is based on the final volume method. The simulations are performed with the spatial resolution, $500\times 100\times 100$ points in $x$, $y$ and $z$ directions, respectively. In the mean-field numerical simulations, we use the following values of the basic dimensionless parameters: the turbulent Prandtl number ${\rm Pr}_{{}_{T}}=1$, the ratio $u_{c}/u_{0}=1$, the effective Rayleigh number changies from ${\rm Ra}_{{}_{T}}=0.5$ to ${\rm Ra}_{{}_{T}}=1800$ and the scale separation parameter $\epsilon$ varies from $\epsilon=10^{-3}$ to $\epsilon=10^{-2}$. The non-dimensional shear number ${\rm Sh}$ varies from ${\rm Sh}=0.01$ to ${\rm Sh}=1$.
Figure 11: The counter lines of the velocity field (upper panel); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (middle panel); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (bottom panel) at time instant $t=0.34$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the stress-free vertical boundary conditions. All quantities are normalized by their maximum values.
@@
For illustration, in Figs. 5 and 6 we plot the time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for different values of the effective Rayleigh numbers ${\rm Ra}_{{}_{T}}$ changing from $0.5$ to $10^{3}$ at a fixed value of the parameter $\epsilon$ in the sheared large-scale convection (for ${\rm Sh}=1$), while in Figs. 710 we show the time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for different values of the scale separation parameter $\epsilon$ (from $10^{-3}$ to $10^{-2}$) between the vertical size $L_{z}$ of the computational domain and the integral turbulence scale $\ell_{0}$ (see Figs. 78 for ${\rm Ra}_{{}_{T}}=0.5$ and Figs. 910 for ${\rm Ra}_{{}_{T}}=10^{3}$).
@@
As can be seen in Figs. 510, at the initial stage of the evolution, the maximum velocity $\tilde{U}_{\rm max}(t)$ increases in time exponentially due to the excitation of the large-scale convective-shear instability. During the nonlinear stage of the instability, we observe that $\tilde{U}_{\rm max}(t)$ reaches the maximum value which is weakly dependent on the effective Rayleigh number ${\rm Ra}_{{}_{T}}$ for the stress-free and no-slip boundary conditions (see Figs. 56). On the other hand, the maximum value of the function $\tilde{U}_{\rm max}$ at the stationary stage strongly depends on the scale separation parameter $\epsilon$ and on the boundary conditions (see Figs. 78). In particular, increasing the scale separation between the vertical size $L_{z}$ of the computational domain and the integral turbulence scale $\ell_{0}$ (i.e., decreasing the parameter $\epsilon$), we observe that the maximum value of the function $\tilde{U}_{\rm max}$ increases. For the no-slip vertical boundary conditions, the large-scale instability is excited and convective structures are formed when $\epsilon<3\times 10^{-3}$ at ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon<4\times 10^{-3}$ at ${\rm Ra}_{{}_{T}}=500$ and $\epsilon<7\times 10^{-3}$ at ${\rm Ra}_{{}_{T}}=10^{3}$. In addition, the time evolution in the nonlinear stage of the instability for the stress-free and no-slip boundary conditions are different.
@@
To observe spatial structure of the basic characteristics of the sheared large-scale convection, in Figs. 1115 we plot the counter lines of the velocity field (upper panels); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (middle panels) and the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (bottom panels) at several time instants: before the system reaches steady-state (see Figs. 11 and 12) and after the system reaches steady-state (see Fig. 1315).
Figure 12: The counter lines of the velocity field (upper panel); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (middle panel); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (bottom panel) at time instant $t=0.7$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the no-slip vertical boundary conditions. All quantities are normalized by their maximum values.
@@
We remind that the potential temperature is measured in the units of $L_{z}\,N^{2}\,{\rm Pr}_{{}_{T}}/\beta={\rm Ra}_{{}_{T}}\nu_{{}_{T}}^{2}/(\beta L_{z}^{3})$. This is the reason why we show in Figs. 1115 (see middle panels) the pattern of the normalized deviations of the potential temperature $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state. Note also that the total gradient of the potential temperature is the sum of the equilibrium constant gradient of the potential temperature $\nabla_{z}\overline{T}_{\rm eq}$ (negative for a convection) and the gradient of the potential temperature $\nabla_{z}\overline{\Theta}$. Therefore, we show in Figs. 1115 (see bottom panels) the pattern of the normalized total vertical gradient of the mean potential temperature, $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ which characterises the large-scale convection.
@@
At the linear stage of the system evolution, the patterns for the stress-free and no-slip boundary conditions are the same (see Fig. 16). During the evolution, there is a transition from the large-scale circulations at the $yz$ plane seen in the linear stage of the instability (see Fig. 16) to the the large-scale circulations at the $xz$ plane observed during nonlinear stage of the instability (see Figs. 1112). In addition, there is a transition from the two-layer vertical structure of the mean velocity field with two convective roles in the $z$ direction and six roles in the $x$ direction (as can be seen in Fig. 11 for the stress-free boundary conditions) to the one-layer vertical structure with one role in the $z$ direction and two roles in the $x$ direction (see Fig. 13).
On the other hand, comparing Figs. 12 and 14 (which corresponds to the no-slip boundary conditions), we observe that during nonlinear stage of the instability there is a transition from the large-scale circulations with the two-layer vertical structure with two convective roles in the $z$ direction and six roles in the $x$ direction (see Fig. 12) to the one-layer structure with four inclined roles (see Fig. 14). The formation of the inclined roles for the no-slip boundary conditions are already observed starting with $t=0.9$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$ at the effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$. The inclined roles for the no-slip boundary conditions are also observed at the effective Rayleigh number ${\rm Ra}_{{}_{T}}=10^{3}$ (see Fig. 15).
@@
Figure 13: The counter lines of the velocity field (upper panel); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (middle panel); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (bottom panel) at time instant $t=1$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the stress-free vertical boundary conditions. All quantities are normalized by their maximum values.
@@
As has been observed in the shear-free convection [53], the large-scale convective structures are also formed in a sheared convection even at low values of the effective Rayleigh numbers ${\rm Ra}_{{}_{T}}=0.5$ due to the additional terms $\propto\epsilon$ in Eq. (5) for the evolution of the potential temperature (which are caused by the modification of the turbulent heat flux by non-uniform fluid flows). Moreover, in Figs. 1115 (bottom panels), one can see the regions with the positive gradient of the potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$, which are typical for stably stratified turbulence. Such effects have been previously observed in experiments [16, 54], direct numerical simulations [44, 46, 55, 47] of turbulent convection and shear-free mean-field numerical simulations [53]. The formation of the regions with the positive gradient of the potential temperature inside the large-scale circulation can be understood as follows. The total vertical heat flux ${\bm{F}}_{z}^{\rm tot}$ includes three contributions [53]:
@@
In the present study, we also observe the formation the large-scale rolls even below the threshold of the laminar convection (see Figs. 1114 for effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$ which can be compared with Fig. 15 for ${\rm Ra}_{{}_{T}}=1000$). This is because turbulence with non-uniform large-scale flows contributes to the turbulent heat flux. The reasons for this effect are related to production of additional essentially anisotropic velocity fluctuations generated by tangling of the mean-velocity gradients by small-scale turbulent motions due to the influence of the inertial forces during the lifetime of turbulent eddies. These anisotropic velocity fluctuations contributes to the turbulent heat flux. As the result of this effect, there is an excitation of large-scale convective-shear instability, which results in the formation of large-scale semi-organized structures in the form of rolls and generation of convective-shear waves propagating perpendicular to the convective rolls. The life-times and spatial scales of these structures are much larger compared to the largest turbulent time scales. As the result, the evolutionary equation (5) for the potential temperature $\tilde{\Theta}$ contains the new terms proportional to the spatial derivatives of the mean velocity field $\tilde{\bm{U}}$ (see the terms $\propto\epsilon$).
@@
Figure 14: The counter lines of the velocity field (upper panel); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (middle panel); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (bottom panel) at time instant $t=4$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the no-slip vertical boundary conditions. All quantities are normalized by their maximum values.
@@
Figure 15: The counter lines of the velocity field (upper panel); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (middle panel); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (bottom panel) at time instant $t=5.32$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=1000$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the no-slip vertical boundary conditions. All quantities are normalized by their maximum values.
@@
Figure 16: The counter lines of the velocity field (upper panel); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (middle panel); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (bottom panel) at time instant $t=0.04$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the no-slip vertical boundary conditions. All quantities are normalized by their maximum values.

### v2

@@ IV Results of mean-field numerical simulations
In this section, we discuss results of mean-field numerical simulations for three-dimensional problem. We solve numerically Eqs. (4)–(5) for the periodic boundary conditions in the horizontal $xy$ plane. The boundary conditions for the potential temperature in the vertical direction are $\tilde{\Theta}(t,z=0)=\tilde{\Theta}(t,z=1)=0$. We adopt the stress-free and no-slip boundary conditions for the velocity field in the vertical direction (along the $z$ axis). The stress-free boundary conditions imply,
@@
Figure 8: Time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\alpha=0.5$ and shear number ${\rm Sh}=1$, and at different values of the scale separation parameter $\epsilon=10^{-3};5\times 10^{-3};10^{-2}$, for the stress-free boundary conditions.
To solve Eqs. (4)–(5), we use the ANSYS FLUENT code (version 19.2) in the $3D$ box $L_{x}=5L_{z}$ and $L_{y}=L_{z}$, which is based on the finite volume method. The discretization method in the code is the second-order. The simulations are performed with the spatial resolution $500\times 100\times 100$ in $x$, $y$ and $z$ directions, respectively. A sensitivity check has been also made for the spatial resolution $500\times 200\times 200$. For both cases similar results for velocity and potential temperature have been obtained. In all simulations, we use a time step of $10^{-3}$. A sensitivity check has been made for time steps as well. Time steps of $10^{-3}$ and $2\times 10^{-3}$ have been tested and a maximum error of 0.05 % at velocity and potential temperature between the time steps has been obtained. In addition, a convergence error was set to be less than $10^{-6}$.
@@
In the mean-field numerical simulations, we use the following values of the basic dimensionless parameters: the turbulent Prandtl number ${\rm Pr}_{{}_{T}}=1$, the ratio $u_{c}/u_{0}=1$, the effective Rayleigh number changes from ${\rm Ra}_{{}_{T}}=0.5$ to ${\rm Ra}_{{}_{T}}=1800$ and the scale separation parameter $\epsilon$ varies from $\epsilon=10^{-3}$ to $\epsilon=10^{-2}$. The non-dimensional shear number ${\rm Sh}$ varies from ${\rm Sh}=0.01$ to ${\rm Sh}=1$. Note that the measurements of the two-point correlation function of the turbulent velocity field in the laboratory experiments with turbulent convection [25, 26, 54] and DNS [55, 56] show that the integral turbulent scale $\ell_{0}$ varies from $L_{z}/10$ to $L_{z}/5$ which corresponds to $\epsilon$ used in our mean-field simulations.
Figure 9: Time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\alpha=0.5$ and shear number ${\rm Sh}=1$, and at different values of the scale separation parameter $\epsilon=10^{-3};5\times 10^{-3};10^{-2}$, for the no-slip boundary conditions.
@@
For illustration, in Figs. 6 and 7 we plot the time evolution of the maximum value $\tilde{U}_{\rm max}(t)$ of the velocity magnitude $|\tilde{\bm{U}}|$ for different values of the effective Rayleigh numbers ${\rm Ra}_{{}_{T}}$ changing from $0.5$ to $10^{3}$ at a fixed value of the scale separation parameter $\epsilon$ between the vertical size $L_{z}$ of the computational domain and the integral turbulence scale $\ell_{0}$ in the sheared large-scale convection (for ${\rm Sh}=1$), while in Figs. 811 we show the time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for different values of the scale separation parameter $\epsilon$ (from $10^{-3}$ to $10^{-2}$) (see Figs. 89 for ${\rm Ra}_{{}_{T}}=0.5$ and Figs. 1011 for ${\rm Ra}_{{}_{T}}=10^{3}$). As can be seen in Figs. 611, at the initial stage of the evolution, the maximum velocity $\tilde{U}_{\rm max}(t)$ increases in time exponentially due to the excitation of the large-scale convective-shear instability. The growth rate of the instability predicted from the theory [as the solution of the cubic dispersion equation (13)] is in agreement with that obtained in the mean-field numerical simulations.
@@
During the nonlinear stage of the instability, we observe that $\tilde{U}_{\rm max}(t)$ reaches the maximum value which is weakly dependent on the effective Rayleigh number ${\rm Ra}_{{}_{T}}$ for the stress-free and no-slip boundary conditions (see Figs. 67). On the other hand, the maximum value of the function $\tilde{U}_{\rm max}$ at the stationary stage strongly depends on the scale separation parameter $\epsilon$ and on the boundary conditions (see Figs. 89). In particular, increasing the scale separation between the vertical size $L_{z}$ of the computational domain and the integral turbulence scale $\ell_{0}$ (i.e., decreasing the parameter $\epsilon$), the maximum value of the function $\tilde{U}_{\rm max}$ increases. For the no-slip vertical boundary conditions, the large-scale instability is excited and convective structures are formed when $\epsilon<3\times 10^{-3}$ at ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon<4\times 10^{-3}$ at ${\rm Ra}_{{}_{T}}=500$ and $\epsilon<7\times 10^{-3}$ at ${\rm Ra}_{{}_{T}}=10^{3}$. In addition, the time evolution in the nonlinear stage of the instability for the stress-free and no-slip boundary conditions are different.
Figure 10: Time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=1000$, $\alpha=0.5$ and shear number ${\rm Sh}=1$, and at different values of the scale separation parameter $\epsilon=10^{-3};5\times 10^{-3};10^{-2}$, for the stress-free boundary conditions.
@@
To observe spatial structure of the basic characteristics of the sheared large-scale convection, in Figs. 1216 we plot the patterns of velocity vectors $\tilde{\bm{U}}=\tilde{U}_{y}\,{\bm{e}}_{y}+\tilde{U}_{z}\,{\bm{e}}_{z}$ with $\tilde{U}_{x}\to 0$ (panels a); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (panels b) and the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (panels c) at several time instants: before the system reaches steady-state (see Figs. 12 and 13) and after the system reaches steady-state (see Fig. 1416). Here ${\bm{e}}_{x}$, ${\bm{e}}_{y}$ and ${\bm{e}}_{z}$ are the unit vectors directed along the $x$, $y$ and $z$ axes.
Figure 11: Time evolution of the maximum velocity $\tilde{U}_{\rm max}(t)$ for the effective Rayleigh number ${\rm Ra}_{{}_{T}}=1000$, $\alpha=0.5$ and shear number ${\rm Sh}=1$, and at different values of the scale separation parameter $\epsilon=10^{-3};5\times 10^{-3};10^{-2}$, for the no-slip boundary conditions.
@@
We remind that the potential temperature is measured in the units of $L_{z}\,N^{2}\,{\rm Pr}_{{}_{T}}/\beta={\rm Ra}_{{}_{T}}\nu_{{}_{T}}^{2}/(\beta L_{z}^{3})$. This is the reason why we show in Figs. 1216 (see panels b) the pattern of the normalized deviations of the potential temperature $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state. Note also that the total gradient of the potential temperature is the sum of the equilibrium constant gradient of the potential temperature $\nabla_{z}\overline{T}_{\rm eq}$ (negative for a convection) and the gradient of the potential temperature $\nabla_{z}\overline{\Theta}$. Therefore, we show in Figs. 1216 (see panels c) the pattern of the normalized total vertical gradient of the mean potential temperature, $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ which characterises the large-scale convection.
@@
At the linear stage of the system evolution, the patterns for the stress-free and no-slip boundary conditions are the same (see Fig. 17 shown in the $yz$ plane, where in the panel (a) we plot the patterns of velocity vectors $\tilde{\bm{U}}=\tilde{U}_{y}\,{\bm{e}}_{y}+\tilde{U}_{z}\,{\bm{e}}_{z}$ with $\tilde{U}_{x}\to 0$). During the evolution, there is a transition from the large-scale circulations at the $yz$ plane seen in the linear stage of the instability (see Fig. 17) to the the large-scale circulations at the $xz$ plane observed during nonlinear stage of the instability (see Figs. 1213). In addition, there is a transition from the two-layer vertical structure of the mean velocity field with two convective rolls in the $z$ direction and six rolls in the $x$ direction (as can be seen in Fig. 12 for the stress-free boundary conditions) to the one-layer vertical structure with one roll in the $z$ direction and two rolls in the $x$ direction (see Fig. 14).
@@
Figure 12: The patterns of velocity vectors $\tilde{\bm{U}}=\tilde{U}_{x}\,{\bm{e}}_{x}+\tilde{U}_{z}\,{\bm{e}}_{z}$ with $\tilde{U}_{y}\to 0$ (Fig. 12a); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (Fig. 12b); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (Fig. 12c) at time instant $t=0.34$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the stress-free vertical boundary conditions. All quantities are normalized by their maximum values.
On the other hand, comparing Figs. 13 and 15 (which corresponds to the no-slip boundary conditions), we observe that during nonlinear stage of the instability there is a transition from the large-scale circulations with the two-layer vertical structure with two convective rolls in the $z$ direction and six rolls in the $x$ direction (see Fig. 13) to the one-layer structure with four inclined rolls (see Fig. 15). The formation of the inclined rolls for the no-slip boundary conditions are already observed starting with $t=0.9$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$ at the effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$. The inclined rolls for the no-slip boundary conditions are also observed at the effective Rayleigh number ${\rm Ra}_{{}_{T}}=10^{3}$ (see Fig. 16).
@@
As has been observed in the shear-free convection [53], the large-scale convective structures are also formed in a sheared convection even at low values of the effective Rayleigh numbers ${\rm Ra}_{{}_{T}}=0.5$ due to the additional terms proportional to $\epsilon$ in Eq. (5) for the evolution of the potential temperature (which are caused by the modification of the turbulent heat flux by non-uniform fluid flows). Moreover, in Figs. 1216 (panels c), one can see the regions with the positive gradient of the potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$, which are typical for stably stratified turbulence. Such effects have been previously observed in experiments [16, 57], direct numerical simulations [44, 46, 58, 47] of turbulent convection and shear-free mean-field numerical simulations [53]. The formation of the regions with the positive gradient of the potential temperature inside the large-scale circulation can be understood as follows. The total vertical heat flux ${\bm{F}}_{z}^{\rm tot}$ includes three contributions [53]:
@@
Figure 13: The patterns of velocity vectors $\tilde{\bm{U}}=\tilde{U}_{x}\,{\bm{e}}_{x}+\tilde{U}_{z}\,{\bm{e}}_{z}$ with $\tilde{U}_{y}\to 0$ (Fig. 13a); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (Fig. 13b); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (Fig. 13c) at time instant $t=0.7$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the no-slip vertical boundary conditions. All quantities are normalized by their maximum values.
In the present study, we also observe the formation the large-scale rolls even below the threshold of the laminar convection (see Figs. 1215 for effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$ which can be compared with Fig. 16 for ${\rm Ra}_{{}_{T}}=1000$). This is because turbulence with non-uniform large-scale flows contributes to the turbulent heat flux. The reasons for this effect are related to production of additional essentially anisotropic velocity fluctuations generated by tangling of the mean-velocity gradients by small-scale turbulent motions due to the influence of the inertial forces during the lifetime of turbulent eddies. These anisotropic velocity fluctuations contribute to the turbulent heat flux. In particular, anisotropic velocity fluctuations ${\bm{u}}^{\rm anisot}$ are produced by the large-scale shear $\nabla_{j}\overline{U}_{i}$ in the presence of isotropic velocity fluctuations ${\bm{u}}^{\rm isot}$, i.e., $\partial{\bm{u}}^{\rm anisot}/\partial t\propto-({\bm{u}}^{\rm isot}\cdot\mbox{\boldmath$\nabla$})\overline{U}_{i}$. This mechanism is called ”tangling”. Similar mechanism exists in fluid dynamics, when temperature fluctuations $\theta$ are produced by ”tangling” of the gradient of the mean temperature $\nabla_{j}\overline{T}$ by velocity fluctuations ${\bm{u}}$, i.e., $\partial\theta/\partial t\propto-({\bm{u}}\cdot\mbox{\boldmath$\nabla$})\overline{T}$. Moreover, the tangling mechanism also exists in magnetohydrodynamics, when magnetic fluctuations ${\bm{b}}$ are produced by ”tangling” of the mean magnetic field $\overline{\bm{B}}$ by velocity fluctuations, i.e., $\partial{\bm{b}}/\partial t\propto(\overline{\bm{B}}\cdot\mbox{\boldmath$\nabla$}){\bm{u}}$. So the ”tangling” is an universal mechanism of production of anisotropic fluctuations.
@@
Figure 14: The patterns of velocity vectors $\tilde{\bm{U}}=\tilde{U}_{x}\,{\bm{e}}_{x}+\tilde{U}_{z}\,{\bm{e}}_{z}$ with $\tilde{U}_{y}\to 0$ (Fig. 14a); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (Fig. 14b); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (Fig. 14c) at time instant $t=1$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the stress-free vertical boundary conditions. All quantities are normalized by their maximum values.
The modification of the turbulent heat flux by anisotropic velocity fluctuations causes an excitation of large-scale convective-shear instability, which results in the formation of large-scale semi-organized structures in the form of rolls and generation of convective-shear waves propagating perpendicular to the convective rolls. The life-times and spatial scales of these structures are much larger compared to the largest turbulent time scales. As the result, the evolutionary equation (5) for the potential temperature $\tilde{\Theta}$ contains the new terms proportional to the spatial derivatives of the mean velocity field $\tilde{\bm{U}}$ (see the terms proportional to the parameter $\epsilon$).
@@
Figure 15: The patterns of velocity vectors $\tilde{\bm{U}}=\tilde{U}_{x}\,{\bm{e}}_{x}+\tilde{U}_{z}\,{\bm{e}}_{z}$ with $\tilde{U}_{y}\to 0$ (Fig. 15a); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (Fig. 15b); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (Fig. 15c) at time instant $t=4$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the no-slip vertical boundary conditions. All quantities are normalized by their maximum values.
Figure 16: The patterns of velocity vectors $\tilde{\bm{U}}=\tilde{U}_{x}\,{\bm{e}}_{x}+\tilde{U}_{z}\,{\bm{e}}_{z}$ with $\tilde{U}_{y}\to 0$ (Fig. 16a); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (Fig. 16b); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (Fig. 16c) at time instant $t=5.32$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=1000$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the no-slip vertical boundary conditions. All quantities are normalized by their maximum values.
Figure 17: The patterns of velocity vectors $\tilde{\bm{U}}=\tilde{U}_{y}\,{\bm{e}}_{y}+\tilde{U}_{z}\,{\bm{e}}_{z}$ with $\tilde{U}_{x}\to 0$ (Fig. 17a); the patterns of the potential temperature deviations $\tilde{\Theta}\,{\rm Ra}_{{}_{T}}$ from the equilibrium potential temperature in the basic reference state (Fig. 17b); the patterns of the vertical gradient of the mean potential temperature $(\nabla_{z}\tilde{\Theta}-1)\,{\rm Ra}_{{}_{T}}$ (Fig. 17c) at time instant $t=0.04$ of the turbulent viscosity time $L_{z}^{2}/\nu_{{}_{T}}$, effective Rayleigh number ${\rm Ra}_{{}_{T}}=0.5$, $\epsilon=10^{-3}$, $\alpha=0.5$ and shear number ${\rm Sh}=1$ for the stress-free vertical boundary conditions. All quantities are normalized by their maximum values.
[…]